### v2

@@ 4.2 WSI Classification
WSI classification holds significant importance in pathology diagnosis. It plays a crucial role in accurately analyzing and interpreting WSI, enabling pathologists to make informed diagnostic decisions. In CPath, WSI classification typically employs multiple instance learning (MIL) as the underlying methodology. The MIL approach involves the following steps: (1) Non-overlapping tissue patches are cropped from the original WSI, and features are extracted using a feature extractor. (2) A feature aggregator is applied to integrate the patch-level features into a slide-level feature, enabling classification. To preprocess the WSIs, we utilize the pipeline described in the CLAM toolkit [71]. Specifically, we employ the default segmentation configuration of CLAM to extract patches with 512512 pixels at level 0 for all slides. Slides with a limited number of patches are discarded. Once all patches are extracted, we resize them to 224224 pixels. We then utilize foundation models to extract features from the resized patches and save these features for subsequent MIL analysis. There are several MIL methods available, such as Attention-Based Multiple Instance Learning (ABMIL) [43] and TransMIL [78]. After evaluating the performance of different foundation models across various WSI classification tasks, we found that ABMIL consistently achieves the best results, which aligns with the findings from previous studies [33, 34]. Therefore, we adopt ABMIL to evaluate the performance of different foundation models in our experiments. The architecture and training details of ABMIL are presented in Extended Data Table 39.
@@
To evaluate the performance of the MIL model, we assess the balanced accuracy, weighted F1 score, and AUC, which consider the class imbalance present in the dataset.
Our experiments encompass 12 pathology WSI classification tasks, including (1) breast cancer metastasis detection, (2) coarse-grained breast carcinoma subtyping, (3) fine-grained breast carcinoma subtyping, (4) lobular & ductal carcinoma subtyping, (5) ovarian cancer subtyping, (6) renal cell carcinoma (RCC) subtyping based on TCGA, (7) non-small cell lung cancer subtyping based on TCGA, (8) non-small cell lung cancer subtyping based on CPTAC, (9) prostate cancer grade assessment, (10) TP53 mutation prediction on LUAD (TCGA), (11) Brain tumor subtyping, and (12) glioma IDH1 mutation prediction.
The results of our experiments are presented in Extended Data Tables 3 to 14.
In Table 2, we showcase the average performance across all 12 WSI classification tasks, demonstrating that our method achieves the best performance.
TCGA-NSCLC for NSCLC Subtyping (2 classes). The TCGA-NSCLC dataset was adopted for the subtyping of NSCLC, specifically lung adenocarcinoma (LUAD) and lung squamous cell carcinoma (LUSC), obtained from the TCGA database [45].
The dataset consists of 541 LUAD slides and 512 LUSC slides.
For training and evaluation, the TCGA-NSCLC cohort is label-stratified into train-validation-test sets in a ratio of 7:1:2, resulting in 738 slides for training, 105 slides for validation, and 210 slides for testing. The experimental results are presented in Extended Data Table 3.
CPTAC-NSCLC for NSCLC Subtyping (2 classes). This dataset was adopted for the subtyping of NSCLC using lung squamous cell carcinoma (LSCC) and LUAD WSIs sourced from the CPTAC data portal[79].
After preprocessing, four LSCC slides and one LUAD slide are excluded due to a limited number of tissues, resulting in a total of 1,077 LSCC slides and 1,136 LUAD slides available for further analysis.
For training and evaluation, the train-validation-test cohort is label-stratified in a ratio of 7:1:2, resulting in 1549 slides for training, 222 slides for validation, and 442 slides for testing. The experimental results are reported in Extended Data Table 4.
@@
TCGA-RCC for RCC Subtyping (3 classes). This task contains kidney renal papillary cell carcinoma (KIRP), kidney chromophobe (KICH) and kidney renal clear cell carcinoma (KIRC) WSIs from TCGA database [45].
@@
For training and evaluation, we label-stratified the TCGA-RCC cohort into 7:1:2 train-validation-test (656:94:187 slides). The experimental results are reported in Extended Data Table 5.
@@
TCGA-BRCA for Lobular and Ductal Carcinoma Subtyping (2 classes). This dataset is derived from the TCGA-BRCA dataset [45], consisting of 787 invasive ductal carcinoma (IDC) slides and 198 invasive lobular carcinoma (ILC) slides. For training and evaluation, the dataset was label-stratified into train-validation-test folds in a ratio of 7:1:2, resulting in 689 slides for training, 99 slides for validation, and 197 slides for testing. The subtyping results are presented in Extended Data Table 7.
@@
For training and evaluation, a label-stratified train-validation-test split is employed, maintaining a ratio of 7:1:2 based on the fine-grained classes. This partitioning results in 382 slides for training, 54 slides for validation, and 109 slides for testing. The coarse-grained and fine-grained classification results are presented in Extended Data Table 8 and Extended Data Table 9, respectively.
@@
TCGA-LUAD for Lung Adenocarcinoma TP53 Gene Mutation Prediction (2 classes). The LUAD TP53 gene mutation prediction task consists of 469 FFPE H&E-stained WSIs of lung adenocarcinoma sourced from the TCGA database, along with their TP53 gene mutation annotations. The slides without reported TP53 mutation status are excluded from the dataset. WSIs used in this task is classified into 2 classes, namely TP53 Mutant (248 slides), and TP53 Wildtype (221 slides). For training and evaluation, we label-stratified the WSIs into a training-validation-test cohort with a ratio of 7:1:2, including 345 slides for training, 41 slides for validation, and 83 slides for testing. The experimental results for TCGA-LUAD TP53 gene mutation prediction could be found in Extended Data Table 11.
TCGA-GBMLGG for Glioma IDH1 Mutation Prediction (2 classes). This dataset is specifically designed for glioma IDH1 mutation prediction. It comprises 804 FFPE H&E-stained WSIs sourced from the TCGA database, focusing on the subsets TCGA-GBM and TCGA-LGG [45].
Note that cases without reported IDH mutation status are excluded from the dataset.
The slides in this task are classified into two classes: IDH-1 Wildtype (600 slides) and IDH-1 Mutant (204 slides).
For training and evaluation, the train-validation-test cohort is label-stratified in a ratio of 7:1:2, resulting in 563 slides for training, 80 slides for validation, and 161 slides for testing.
The experimental results for this task can be found in Extended Data Table 12.
UBC-OCEAN for Ovarian Cancer Subtyping (5 classes). This dataset is a collection of 538 slides obtained from the UBC Ovarian Cancer subtypE clAssification and outlier detectioN (UBC-OCEAN) competition [49, 80].
@@
After image preprocessing, the slides without sufficient foregrounds are excluded to reduce data noise, resulting in a total of 527 slides for further analysis. The subtypes of the dataset contains: Clear Cell (CC, 98 slides), Endometrioid (EC, 122 slides), High-Grade Serous Carcinoma (HGSC, 221 slides), Low-Grade Serous Carcinoma (LGSC, 43 slides) , and Mucinous Carcinoma (MC, 43 slides).
@@
EBRAINS for Brain Cancer Subtyping (12 classes).
This brain tumor subtyping task contains 2,323 slides from EBRAINS Digital Tumor Atlas [50].
To ensure reliable subtyping results, we have focused on subtypes with a minimum of 26 slides.
Additionally, we employed the OncoTree classification system [81] to merge subtypes and generate a final set of 12 classes. These classes include Metastatic Tumor (46 slides), Diffue Glioma (885 slides), Ependymomal Tumor (92 slides), Soft Tissue (66 slides), Embryonal Tumor (32 slides), Encapsulated Glioma (260 slides), Sellar Tumor (182 slides), Chordoma (28 slides), Meningothelial Tumor (476 slides), Miscellaneous Brain Tumor (146), Schwannoma (81 slides), Langerhans Cell Histiocytosis (29 slides).
For training and evaluation, we label-stratified the train-validation-test cohort into 7:1:2 (1627:233:463 slides).
The experimental results for this task are reported in Extended Data Table 14.
@@ 4.3 Survival Analysis
Survival analysis has traditionally been employed to analyze time-to-event data in cancer studies, focusing on events such as disease progression or patient survival. When applied to WSIs, survival analysis offers new opportunities for studying various aspects of tissue behavior and predicting patient outcomes [55, 82]. By integrating survival analysis with WSIs, researchers can investigate the correlation between specific morphological features and patient outcomes. In our study, we adopt ABMIL [43] for survival analysis with Negative Log-Likelihood (NLL) loss [83], following a similar model architecture and training configuration as WSI classification reported in Extended Data Table 39.
To evaluate the effectiveness of different foundation models in survival analysis, we employ a 5-fold cross-validation setting and utilize the C-index metric to assess performance.
We conduct survival analysis on 11 TCGA datasets, including TCGA-BRCA, TCGA-BLCA, TCGA-KIRC, TCGA-KIRP, TCGA-LUAD, TCGA-STAD, TCGA-LUSC, TCGA-COADREAD, TCGA-GBMLGG, TCGA-SKCM, TCGA-CESC, and perform external validation on CPTAC-LUAD dataset.
The experimental results are presented in Extended Data Table 16 - Table 19.
Additionally, Table 15 demonstrates the average performance across all 12 survival analysis tasks, highlighting that our method achieves the highest performance in terms of the C-index metric.
Breast Cancer Survival Prediction. We leverage the TCGA-BRCA dataset for breast cancer survival prediction, which consists of 1,023 cases (1,089 WSIs).
To ensure robust and consistent results, we maintain uniform censorship (survival status information) between the training and testing datasets.
To address the challenge of imbalanced survival times, we employ a stratified approach.
Specifically, we sort the cases based on survival time and divide them into four equally sized bins.
We assign the label of the bin to all cases within it.
As a result, for each fold of the 5-fold cross-validation, we label-stratify the train-test cohort into an 8:2 ratio, with 821 cases assigned to the training set and 202 cases allocated to the test set.
The mean C-Index and standard deviation obtained from the 5-fold cross-validation results are presented in Extended Data Table 16.
Bladder Cancer Survival Prediction. For bladder cancer survival prediction, we utilize the TCGA-BLCA dataset, which includes 376 patients and a total of 446 slides.
Following a similar approach used in the BRCA task, we employ 5-fold cross-validation to assess the performance of different foundation models.
In each fold of the cross-validation, we stratify the train-test cohort using an 8:2 ratio, with 305 cases assigned to the training set and 71 cases allocated to the test set.
This label-stratified approach guarantees a comprehensive assessment of the models, ensuring their performance is representative.
The results obtained for bladder cancer survival prediction are presented in the BLCA column of Extended Data Table 16.
Kidney Renal Clear Cell Carcinoma Survival Prediction. In this task, we utilized the TCGA-KIRC dataset, which encompasses a cohort of 498 patients, accompanied by a total of 504 slides. By following a similar approach to the BRCA dataset processing, we employed a 5-fold cross-validation methodology to facilitate accurate survival prediction.
To ensure the reliability and robustness of our analysis, we implemented a label-stratified train-test cohort division for each fold, maintaining an 8:2 ratio. Consequently, the training set comprised 401 cases, while the testing set consisted of 97 cases, enabling a comprehensive evaluation survival prediction.
The experimental results are presented in Extended Data Table 16.
Kidney Renal Papillary Cell Carcinoma Survival Prediction. To enable accurate survival prediction for KIRP, we utilized a dataset consisting of 261 patients (285 slides) obtained from TCGA-KIRP.
The slides are preprocessed similar to that employed for the BRCA survival analysis task. To ensure robustness in our survival prediction results, we employed a 5-fold cross-validation approach.
For each fold, we carefully stratified the train-test cohort, ensuring a balanced representation with an 8:2 ratio. This resulted in a training set of 211 cases, while 50 cases were allocated for testing purposes. The experimental results are presented in the KIRP column of Extended Data Table 17.
Stomach Adenocarcinoma Survival Prediction. For accurate survival prediction in stomach adenocarcinoma, we utilized a dataset obtained from TCGA-STAD, which consisted of 363 patients along with 389 slides. Similar to the approach used in the BRCA analysis, we employed a 5-fold cross-validation methodology to evaluate the performance of different foundation models.
For each fold, we meticulously stratified the train-test cohort, ensuring a balanced distribution with a ratio of 8:2 (293:70 cases). This resulted in a training set of 293 cases, while 70 cases were reserved for testing purposes. The detailed experimental results are presented in the STAD column of Extended Data Table 17.
Cervical Squamous Cell Carcinoma and Endocervical Adenocarcinoma Survival Prediction. To provide effective and accurate survival analysis results for patients with cervical squamous cell carcinoma and endocervical adenocarcinoma (CESC), we leveraged a dataset consisting of 250 cases (260 slides in total) obtained from TCGA-CESC. Utilizing the same experimental settings in the BRCA analysis, a 5-fold cross-validation methodology is applied to evaluate the performance of different foundation models.
For each fold, we meticulously stratified the train-test cohort, resulting in a balanced distribution with a ratio of 8:2, including 203 cases for training and 47 cases for testing. The detailed experimental results are shown in the CESC column of Extended Data Table 17.
Lung Adenocarcinoma Survival Prediction. To facilitate accurate survival prediction for lung adenocarcinoma, we utilized a dataset obtained from TCGA-LUAD, consisting of 455 cases (518 slides).
Following a similar processing approach to that employed for BRCA analysis, we also employed a 5-fold cross-validation strategy for LUAD evaluation.
During each fold, we meticulously label-stratified the train-test cohort with an 8:2 ratio, resulting in 366 cases for training and 89 cases for testing.
The mean and standard deviation of the C-Index across the 5 folds are presented in the LUAD column of Extended Data Table 18.
Additionally, we conducted external validation by collecting an additional dataset of 102 lung adenocarcinoma cases (comprising 478 slides) from CPTAC. To perform the external validation, we utilized the model trained on TCGA-LUAD cohorts and directly conducted inference on the CPTAC-LUAD cases.
We evaluated the results for all 5 folds and calculated the mean and standard deviation of the C-Index.
The experimental results are presented in the LUAD-CPTAC column of Extended Data Table 18.
Lung Squamous Cell Carcinoma Survival Prediction. In this study, we utilized a dataset sourced from TCGA, comprising 452 cases (484 slides), for the purpose of survival prediction in lung squamous cell carcinoma (LUSC).
Similar to the data processing methodology employed for BRCA survival prediction, we conducted a 5-fold cross-validation approach to assess the performance of various foundation models.
During each fold, we meticulously label-stratified the train-test cohort with a ratio of 8:2, resulting in 365 cases for training and 87 cases for testing.
The mean and standard deviation of the C-Index across the 5 folds are reported in the LUSC column of Extended Data Table 18.
Colon & Rectum Adenocarcinoma Survival Prediction. In this survival prediction task, we utilized a combined dataset comprising 426 cases (434 slides) from TCGA-COAD and 153 cases (154 slides) from TCGA-READ. By merging these two datasets, we obtained a larger dataset consisting of 579 cases (588 slides).
Following a similar processing approach to that employed for BRCA, we conducted a 5-fold cross-validation to evaluate the performance of different foundation models.
During each fold, we label-stratified the train-test cohort with a ratio of 8:2, thereby resulting in 464 cases for training and 115 cases for testing. The mean and standard deviation of the C-Index across the 5 folds are reported in Extended Data Table 19.
Glioma Survival Prediction. In this glioma survival prediction task, we utilized a comprehensive dataset comprising 370 cases (823 slides) of glioblastoma multiforme (GBM) sourced from TCGA-GBM, and 460 cases (778 slides) of low-grade glioma (LGG) sourced from TCGA-LGG.
By merging these two datasets, we created a larger dataset consisting of 830 cases (1,601 slides) for glioma survival prediction analysis.
Following a data processing approach similar to that employed for BRCA, we performed a 5-fold cross-validation to evaluate the performance of the models. During each fold, we carefully label-stratified the train-test cohort with a ratio of 8:2, resulting in 667 cases for training and 163 cases for testing.
The mean and standard deviation of the C-Index values across the 5 folds are reported in Extended Data Table 19.
Skin Cutaneous Melanoma Survival Prediction. In this task, we focus on predicting survival outcomes in skin cutaneous melanoma.
The dataset used for this task consists of 415 cases, comprising 456 slides, sourced from the TCGA-SKCM.
Similar to the data processing methodology employed in BRCA survival prediction, we employed a 5-fold cross-validation approach to evaluate the performance of different foundation models.
During each fold of the cross-validation process, we meticulously stratified the train-test cohort, ensuring a ratio of 8:2 between the training and testing sets. This resulted in 337 cases being allocated for training purposes, while 78 cases were reserved for testing.
The mean and standard deviation of the C-Index across the 5 folds are reported in the SKCM column of Extended Data Table 19.

### v3

@@ 4.2 WSI Classification
In CPath, WSI classification typically employs multiple instance learning (MIL) as the underlying methodology. The MIL approach involves the following steps: (1) Non-overlapping tissue patches are cropped from the original WSI, and features are extracted using a feature extractor. (2) A feature aggregator is applied to integrate the patch-level features into a slide-level feature, enabling classification. To preprocess the WSIs, we utilize the pipeline described in the CLAM toolkit [58]. Specifically, we employ the default segmentation configuration of CLAM to extract patches with 512512 pixels at level 0 for all slides. Slides with a limited number of patches are discarded. Once all patches are extracted, we resize them to 224224 pixels. We then utilize FMs to extract features from the resized patches and save these features for subsequent MIL analysis. There are several MIL methods available, such as Attention-Based Multiple Instance Learning (ABMIL) [67] and TransMIL [68]. After evaluating the performance of different FMs across various WSI classification tasks, we found that ABMIL consistently achieves the best results, which aligns with the findings from previous studies [33, 34]. Therefore, we adopt ABMIL to evaluate the performance of different FMs in our experiments. The architecture and training details of ABMIL are presented in Extended Data Table 47. For CHIEF [64] and Prov-Gigapath [56] models, we use their pretrained slide-level FM to perform classification.
@@
To evaluate the performance of the MIL model, we assess the balanced accuracy, weighted F1 score, and AUC, which consider the class imbalance present in the dataset. Our experiments encompass 36 pathology WSI classification tasks, including 20 internal and 16 external validation datasets.
The results of our experiments are presented in Extended Data Tables 1-15.
NSCLC Subtyping on TCGA, CPTAC and Center-1 Cohorts (2 classes).
To perform subtyping of non-small cell lung cancer (NSCLC), we utilized data from the TCGA [69], CPTAC [70], and Center-1.
The TCGA cohort comprises 541 lung adenocarcinoma (LUAD) and 512 lung squamous cell carcinoma (LUSC) samples. The data is label-stratified in a ratio of 7:1:2, resulting in 738 slides for training, 105 slides for validation, and 210 slides for testing.
For the CPTAC cohort, there are 1,077 LUSC slides and 1,136 LUAD slides. Similarly, this cohort is label-stratified in a 7:1:2 ratio, yielding 1,549 slides for training, 222 slides for validation, and 442 slides for testing. Additionally, we included 180 LUAD slides and 30 LUSC slides from Center-1 for external validation.
We directly predicted the subtype of the slides using the model trained on the TCGA cohort. The experimental results are presented in Extended Data Table 2.
Lung Cancer Metastatic Detection and Primary Site Prediction (2 classes and 6 classes).
For metastatic detection, we utilized 1,198 WSIs from the Center-1, comprising 705 patients, including 391 primary cases and 314 metastatic cases. To predict the primary site of metastatic cancer, we curated a dataset with six distinct classes: LUAD (391 cases), breast (55 cases), colon (186 cases), kidney (25 cases), liver (34 cases), and carcinoma of unknown primary (CUP, 14 cases). For both tasks, the data were stratified into training, validation, and test sets at a ratio of 7:1:2.
Additionally, we incorporated an external validation cohort consisting of 530 WSIs (431 cases) from Center-2. For the metastatic detection task, the Center-2 cohort included 238 primary cases and 193 metastatic cases. For the primary site prediction task, the Center-2 cohort comprised 238 LUAD cases, 50 breast cases, 96 colon cases, 30 kidney cases, 10 liver cases, and 7 CUP cases.
To facilitate distinction between the datasets, we designated the Center-1 cohort as Center-1-LMD and the Center-2 cohort as Center-2-LMD. The experimental results are presented in Extended Data Table 3.
@@
RCC Subtyping (3 classes) on TCGA and Center-3 Cohorts. This task contains kidney renal papillary cell carcinoma (KIRP), kidney chromophobe (KICH) and kidney renal clear cell carcinoma (KIRC) WSIs from TCGA database [69].
@@
For training and evaluation, we label-stratified the TCGA-RCC cohort into 7:1:2 train-validation-test (656:94:187 slides).
Additionally, we adopted 28 KICH slides, 30 KIRC slides, and 30 KIRP slides from Center-3 (Center-3-RCC) as the external cohort.
The experimental results are reported in Extended Data Table 4.
@@
Lobular and Ductal Carcinoma Subtyping on TCGA and Center-3 Cohorts (2 classes).
We utilized the TCGA-BRCA dataset [69] and slides from the Center-3 for both internal and external experiments. The TCGA-BRCA dataset contains 787 slides of invasive ductal carcinoma (IDC) and 198 slides of invasive lobular carcinoma (ILC). For training and evaluation, the dataset was stratified by labels into training, validation, and testing folds in a ratio of 7:1:2, resulting in 689 slides for training, 99 slides for validation, and 197 slides for testing.
We also adopted BRCA slides (Center-3-LD) from Center-3 to conducte external validation.
This dataset comprises 84 ILC slides and 299 IDC slides.
The subtyping results are presented in Extended Data Table 6.
@@
For training and evaluation, a label-stratified train-validation-test split is employed, maintaining a ratio of 7:1:2 based on the fine-grained classes. This partitioning results in 382 slides for training, 54 slides for validation, and 109 slides for testing.
Additionally, we also adopted 84 normal slides and 383 abnormal slides from Center-3 to perform external validation (Center-3-BRCA).
The coarse-grained and fine-grained classification results are presented in Extended Data Table 7 and 8, respectively.
@@
TCGA-LUAD for Lung Adenocarcinoma TP53 Gene Mutation Prediction (2 classes). The LUAD TP53 gene mutation prediction task consists of 469 FFPE H&E-stained WSIs of lung adenocarcinoma sourced from the TCGA database, along with their TP53 gene mutation annotations. The slides without reported TP53 mutation status are excluded from the dataset. WSIs used in this task are classified into 2 classes, namely TP53 Mutant (248 slides), and TP53 Wildtype (221 slides). For training and evaluation, we label-stratified the WSIs into a training-validation-test cohort with a ratio of 7:1:2, including 345 slides for training, 41 slides for validation, and 83 slides for testing. The experimental results for TCGA-LUAD TP53 gene mutation prediction could be found in Extended Data Table 10.
The mutation Status of IDH in Glioma (2 classes).
To predict the IDH mutational status in gliomas, we utilized data from TCGA-GBM and TCGA-LGG, comprising a total of 979 slides, including 722 positive slides and 257 negative slides. For model training and evaluation, the dataset was divided into training, validation, and test sets in a label-stratified ratio of 7:1:2. Additionally, to validate the robustness of our model, we incorporated an external validation set consisting of 852 slides (322 positives and 530 negatives) from EBRAINS [75]. The detailed experimental results for this task are presented in Extended Data Table 11.
Ovarian Cancer Subtyping (5 classes) on UBC-OCEAN and Center-3 Cohorts.
To perform overian cancer classification, we adopted UBC-OCEAN dataset. This dataset is a collection of 538 slides obtained from the UBC Ovarian Cancer subtypE clAssification and outlier detectioN (UBC-OCEAN) competition [76, 77].
@@
After image preprocessing, the slides without sufficient foregrounds are excluded to reduce data noise, resulting in a total of 527 slides for further analysis. The subtypes of the dataset contains Clear Cell (CC, 98 slides), Endometrioid (EC, 122 slides), High-Grade Serous Carcinoma (HGSC, 221 slides), Low-Grade Serous Carcinoma (LGSC, 43 slides) , and Mucinous Carcinoma (MC, 43 slides).
@@
In addition, we also adopted 100 CC, 100 HGSC, 38 LGSC, 97 EC and 35 MC slides from Center-3 as the external validation cohort (Center-3-Ovary).
@@
Brain Tumor Subtyping (3 classes).
To conduct brain tumor subtyping, we utilized a dataset of 1,276 slides from TCGA-GBM and TCGA-LGG, comprising 217 oligodendroglioma slides, 164 anaplastic astrocytoma slides, and 895 glioblastoma slides. For model training and evaluation, the dataset was label-stratified and divided into training, validation, and test sets with 839, 200, and 237 slides, respectively. Additionally, we incorporated an external validation set of 732 slides from the EBRAINS Digital Tumor Atlas [75], which includes 84 oligodendroglioma slides, 89 anaplastic astrocytoma slides, and 559 glioblastoma slides. The experimental results for this task are detailed in Extended Data Table 13.
Lesion grade Classification of Colon Cancer.
To perform lesion grade classification in colon cancer, we utilized the IMP-CRS-2024 dataset [78, 79, 80] for experiments.
This dataset comprises 847 non-neoplastic slides, 2,847 low-grade lesion slides, and 1,638 high-grade lesion slides.
We adhered to the official dataset splits, using 3,300 slides from CRS2 for training, 1,132 slides from CRS1 for validation, and 900 slides from CRS_Test for testing. Additionally, we incorporated an external validation set from Center-3, referred to as Center-3-Colon-WSI, which includes 100 non-neoplastic slides, 121 low-grade lesion slides, and 76 high-grade lesion slides. The experimental results for this task are detailed in Extended Data Table 14.
Head&Neck Cancer Primary Site Prediction and TNM analysis
We employed the HANCOCK dataset [81] to predict the primary site of head and neck tumors and to determine the T stage of the tumors.
For primary site prediction, we utilized 708 slides, including 80 hypopharynx slides, 182 larynx slides, 317 oropharynx slides, and 129 oral cavity slides. The dataset was label-stratified and divided into 495 WSIs for training, 68 WSIs for validation, and 145 WSIs for testing.
For the TNM analysis task, we used 705 slides from the HANCOCK dataset to predict the tumor stage (T stage). This dataset comprises 259 T1 slides, 256 T2 slides, 123 T3 slides, and 67 T4 slides. The dataset was partitioned into training, validation, and testing sets with 496, 67, and 142 slides, respectively.
The experimental results for both tasks are presented in Extended Data Table 15.
Lauren Subtyping of Gastric Cancer.
We utilized the TCGA-STAD dataset to conduct Lauren classification. The TCGA-STAD cohort comprises 81 diffuse-type, 125 mixed-type, and 184 intestinal-type WSIs.
For model training and evaluation, we divided the dataset into training, validation, and test sets in a stratified 7:1:2 ratio based on labels. Furthermore, we incorporated 141 WSIs from the Center-5 and 319 WSIs from Center-4 as external validation cohorts. The Center-5 cohort consists of 77 diffuse-type, 33 mixed-type, and 31 intestinal-type WSIs, while the Center-4 cohort includes 143 diffuse-type, 86 mixed-type, and 90 intestinal-type WSIs. We detail the results of these three datasets for this task in Extended Data Table 16.
Vascular Invasion Detection in Gastric Cancer.
To detect vascular invasion in gastric cancer, we utilized a dataset comprising 396 WSIs from Center-1, referred to as the Center-1-Vascular dataset. This dataset includes 197 positive cases and 168 negative cases. For the purpose of model training and evaluation, the data was partitioned into training, validation, and test sets in a ratio of 7:1:2. Additionally, we incorporated two external validation sets: 230 WSIs (140 positive and 90 negative) from Center-5 and 319 WSIs (122 positive and 197 negative) from Center-4. The experiment results of all three datasets of this task are shown in Extended Data Table 17.
Perineural Invasion Detection in Gastric Cancer.
To detect perineural invasion in gastric cancer, we utilized a dataset consisting of 397 WSIs obtained from Center-1.
This dataset includes 255 positive cases and 141 negative cases. For model training and evaluation, the data was divided into training, validation, and test sets in a ratio of 7:1:2.
Furthermore, we incorporated two additional external validation sets: 232 WSIs (156 positive and 76 negative) from Center-5 and 319 WSIs (112 positive and 207 negative) from Center-4. See Extended Data Table 18 for experimental results.
@@ 4.3 Survival Analysis
Survival analysis has traditionally been employed to analyze time-to-event data in cancer studies, focusing on events such as disease progression or patient survival. When applied to WSIs, survival analysis offers new opportunities for studying various aspects of tissue behavior and predicting patient outcomes [47, 82]. By integrating survival analysis with WSIs, researchers can investigate the correlation between specific morphological features and patient outcomes. In our study, we adopt ABMIL [67] for survival analysis with Negative Log-Likelihood (NLL) loss [83], following a similar model architecture and training configuration as WSI classification reported in Extended Data Table 47. For CHIEF and Prov-Gigapath models, we use their pretrained slide-level FM to perform classification.
To evaluate the effectiveness of different FMs in survival analysis, we employ a train:test split of 8:2 setting and utilize the C-index metric to assess performance. We conduct survival analysis on 14 TCGA datasets, including breast cancer (BRCA), bladder cancer (BLCA), kidney renal clear cell carcinoma (KIRC), kidney renal papillary cell carcinoma (KIRP), lung adenocarcinoma (LUAD), stomach adenocarcinoma (STAD), lung squamous cell carcinoma (LUSC), colon adenocarcinoma (COAD), rectum adenocarcinoma (READ), glioblastoma multiforme (GBM), low-grade glioma (LGG), skin cutaneous melanoma (SKCM), cervical squamous cell carcinoma (CESC), and head-neck squamous cell carcinoma (HNSC). Additionally, we performed external validation on the HANCOCK dataset. The number of slides for each dataset is reported in the Extended Data Table 48. To ensure robust and consistent results, we maintain uniform censorship (survival status information) between the training and testing datasets. To address the challenge of imbalanced survival times, we employ a stratified approach. Specifically, we sort the cases based on survival time and divide them into four equally sized bins. We assign the label of the bin to all cases within it. As a result, we label-stratify the train-test cohort into an 8:2 ratio. The experimental results are presented in Extended Data Table 19-23.